A call option is currently selling for $5.30. It has a strike price of $60 and six months to maturity. A put option with the same strike price sells for $7.80. The risk-free rate is 4.3 percent, and the stock will pay a dividend of $2.80 in three months. What is the current stock price?
Answer #1


Using put call parity
S+P=C+X/(1+r)^t+PV(D)
=>S=C+X/(1+r)^t+PV(D)-P
=>S=5.30+60/(1+4.3%)^(6/12)+2.80/(1+4.3%)^(3/12)-7.80
=>S=59.02084947